# Beyond the Doors of Perception: Vision Transformers Represent Relations Between Objects

### v1

@@ Appendix B ViT B/16 All Model Results
See Table 2 for behavioral results from all models trained on discrimination and RMTS tasks.
Pretrain Task Train Acc. IID Test Acc.
CLIP Disc. 100% 99.3%
CLIP RMTS 100% 98.3%
ImageNet Disc. 100% 97.5%
ImageNet RMTS 99.7%89.3%
DINO Disc. 100% 95.6%
DINO RMTS 100% 87.7%
Disc. 95.9%80.5%
RMTS 54.7% 49.6%
@@
Table 2: All behavioral results for ViT-B/16 models.
@@ C.1 Attention Pattern Analysis
Figure 8: CLIP B/32 attention pattern analysis.
@@ C.3 Relational Stage Analysis
See Figure 10.
Figure 10: CLIP B/32 relational stage analysis.
@@ C.4 Generalization Results
Appendix D Attention Patterns Examples
See Figure 11.
Figure 11: Example attention head patterns for CLIP ViT-B/16 and From Scratch ViT-B/16 trained on the Discrimination task.
Appendix E Distributed Alignment Search Technical Details
@@
Appendix F Perceptual Stage Analysis Controls
@@
Appendix G Perceptual Stage Analysis: Other Models
See Figures 14,15,16 for DINO, ImageNet, and From Scratch DAS results.
@@ Appendix G Perceptual Stage Analysis: Other Models
Appendix H Novel Representations Analysis Technical Details
@@
Appendix I Relational Stage Analysis: Other Models
I.1 Novel Representation Analysis
See Figures 18,19,20,21,22 for DINO Discrimination/RMTS, ImageNet Discrimination/RMTS, and From Scratch Discrimination Novel Representation Analysis results.
@@ I.1 Novel Representation Analysis
I.2 Abstract Representations of Same and Different
We run the linear probe and linear intervention analysis from Section 5 on DINO B/16 and ImageNet B/16. We find that the intervention works much less well on these models than on CLIP B/16 or B/32. This indicates that these models are not using one abstract representation of same and one representation of different that is agnostic to perceptual qualities of the input image.
Additionally, we try to scale the directions that we are adding to the intermediate representations by 0.5 and 2.0, and find that neither of these versions of the intervention work much better well for either model. See Figure 23 and 24 for DINO and ImageNet results.
@@ I.2 Abstract Representations of Same and Different
Appendix J Relational Failure Mode
Adding an auxiliary loss term to encourage disentangled object representations does not help for RMTS (See 4).
Aux. Loss Task Patch Size Train Acc. Test Acc. Comp. Acc.
RMTS 16 49.2% 50.1% 50.0%
RMTS 16 53.9% 54.4% 54.1%
RMTS 32 87.7% 51.4% 51.4%
RMTS 32 70.2% 50.2% 50.2%
Table 4: Adding an auxiliary loss term to encourage disentangled object representations does not help for RMTS.
Appendix K Compute Resources
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section 8.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: [N/A]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Sections 2,3,4, 5, Appendices EH
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We link to an anonymized repository containing our code.
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  25. 6.
    Experimental Setting/Details
  26. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  27. Answer: [Yes]
  28. Justification: Sections 2,3,4, 5, Appendices EH
  29. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  30. 7.
    Experiment Statistical Significance
  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  32. Answer: [No]
  33. Justification: Our results primarily concern analyses conducted on finetuned versions of pretrained models. We present model analyses separately so as not to lose information through aggregation, at the expense of reporting statistical significance.
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  35. 8.
    Experiments Compute Resources
  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  37. Answer: [Yes]
  38. Justification: See Appendix K.
  39. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  40. 9.
    Code Of Ethics
  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  42. Answer: [Yes]
  43. Justification: [N/A]
  44. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  45. 10.
    Broader Impacts
  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  47. Answer: [N/A]
  48. Justification: This work centers around issues at the intersection of cognitive science and deep learning. This work does not contribute toward practical algorithms or techniques that will see broad deployment.
  49. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  50. 11.
    Safeguards
  51. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  52. Answer: [N/A]
  53. Justification: The models used in this paper are standard vision transformers finetuned on a simple synthetic task. There are no such risks.
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: [N/A]
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New Assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [N/A]
  63. Justification: [N/A]
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and Research with Human Subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification: [N/A]
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Justification: [N/A]
  74. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@ Appendix B ViT B/16: All Model Behavioral Results
Table 2: All behavioral results for ViT-B/16 models trained on all 256 shape-color combinations on the discrimination task.
Pretrain Pretraining Scale $\downarrow$ Train Acc. Test Acc. Comp. Acc.
CLIP 400M 98.1% 98.5% 98.5%
DINOv2 142M 99.6% 98.5% 98.5%
ImageNet 14.2M 98.1% 95.7% 95.7%
DINO 1.28M 98.1% 92.3% 94.7%
MAE 1.28M 98.1% 94.9% 94.9%
77.3% 76.5% 76.0%
@@
Table 3: All behavioral results for ViT-B/16 models trained on 32 shape-color combinations on the discrimination task.
Pretrain Pretraining Scale $\downarrow$ Train Acc. Test Acc.
CLIP 400M 100% 98.3%
DINOv2 142M 100% 98.2%
ImageNet 14.2M 99.7% 89.3%
DINO 1.28M 100% 87.7%
MAE 1.28M 100% 93.4%
49.2% 50.1%
Table 4: All behavioral results for ViT-B/16 models trained on all 256 shape-color combinations on the RMTS task.
Pretrain Pretraining Scale $\downarrow$ Train Acc. Test Acc. Comp. Acc.
CLIP 400M 100% 98.0% 98.3%
DINOv2 142M 100% 96.4% 94.7%
ImageNet 14.2M 99.5% 92.3% 84.0%
DINO 1.28M 99.6% 94.7% 85.2%
MAE 1.28M 99.6% 85.3% 85.3%
49.2% 50.0% 50.0%
Table 5: All behavioral results for ViT-B/16 models trained on 32 shape-color combinations on the RMTS task.
@@ C.1 Attention Pattern Analysis
Figure 11: CLIP B/32 attention pattern analysis. See the caption of Figure 2 for figure and legend descriptions. The B/32 model follows the same stages of processing as CLIP ViT-B/16, and WO & WP peak at the same layers (3 and 6 for discrimination respectively; 3 and 5 for RMTS respectively). However, WO attention remains high for longer than B/16 models.
@@ C.3 Relational Stage Analysis
See Figure 13 for novel representations analysis. See Figure 14 for linear intervention analysis. We find broadly similar results as CLIP B/16.
Figure 12: CLIP B/32 DAS results.
Figure 13: CLIP B/32 relational stage analysis: Novel Representations.
Figure 14: CLIP B/32 relational stage analysis: Linear Intervention.
@@ C.4 Generalization Results
Appendix D DINOv2 Analyses
D.1 Attention Map Analysis
Figure 15: DINOv2 attention pattern analysis. See the caption of Figure 2 for figure and legend descriptions. Note that the stars in the line charts are placed differently in this figure compared to other attention pattern analysis figures. Instead of marking the maximal values of each type of attention across all 12 layers, the stars mark the maximal value excluding the 0th layer. This is because all types of attention spike in DINOv2 in the 0th layer.
See Figure 15 for DINOv2 attention pattern analyses. Like CLIP, DINOv2 displays two-stage processing (albeit somewhat less cleanly). One notable difference compared to CLIP is that all types of attention (WO, WP, BP, and BG) spike in the 0th layer. This might be related to DINOv2’s positional encodings. Since the model was pretrained on images with a size of $518\times 518$ pixels, the model’s positional encodings are interpolated to process our $224\times 224$ stimuli; this might cause an artifact in the attention patterns in the very beginning of the model. Disregarding this spike, the stages of processing follow CLIP. In the discrimination task (Figure 15a), within-object attention peaks at layer 3 (disregarding the initial peak), followed by within-pair and finally background attention. In the RMTS task (Figure 15b), within-object attention peaks at layer 3, followed by within-pair attention at layer 8, and finally between-pair attention in the final layer. Background attention remains relatively high throughout the model, indicating that DINOv2 might make greater use of register tokens to solve the RMTS task compared to other models.
D.2 Perceptual Stage Analysis
See Figure 16 for perceptual stage analysis of DINOV2-pretrained ViTs. Overall, we find highly disentangled object representations in these models.
Figure 16: DAS results for DINOv2 ViT-B/14.
D.3 Relational Stage Analysis
Figure 17: Novel Representation Analysis for DINOv2 ViT-B/14 (Disc.).
Figure 18: Novel Representation Analysis for DINOv2 ViT-B/14 (RMTS).
D.3.1 Novel Representation Analysis
See Figure 17 and  18 for novel representation analysis of DINOV2-pretrained ViTs for the discrimination and RMTS tasks. These results replicate those found using CLIP-pretrained ViTs.
D.3.2 Abstract Representations of Same and Different
See Figure 19 for linear probing and intervention results for DINOV2-pretrained ViTs. We find that the intervention works extremely well for these models, replicating our results on CLIP-pretrained ViTs.
Figure 19: Linear probe & intervention analysis for DINOv2 ViT-B/14.
Appendix E Attention Pattern Analyses
Figure 20: ImageNet, DINO, and MAE ViT attention pattern analysis. See the caption of Figure 2 for figure and legend descriptions. Like CLIP and DINOv2, ImageNet ViT displays two-stage processing on both the discrimination and RMTS tasks; however, performance of this model lags behind CLIP and DINOv2, possibly due to smaller pretraining scale (see the Pretraining Scale column in Table 2). DINO and MAE do not display two-stage processing. These two models are also pretrained on the smallest amount of data, further supporting our intuition that pretraining scale rather than objective results in two-stage processing.
See Figure 20 for attention pattern analyses for ImageNet, DINO, and MAE ViT on the Discrimination and RMTS tasks. ImageNet loosely demonstrates two-stage processing like CLIP and DINOv2. On the other hand, DINO and MAE do not display two stage processing; instead, local and global processing appears to be mixed throughout the models. DINO and MAE also receive the smallest scale pretraining compared to the other models (see the Pretraining Scale column in Table 2); this provides further support for our intuition that pretraining scale results in two-stage processing.
Appendix F Two Internal Algorithms Examined in Greater Detail
While the attention head analysis in Section 3 shows that different models use qualitatively different internal algorithms to solve same-different tasks, the specific computations involved in these algorithms are less clear. What exactly is happening during CLIP’s perceptual stage, for example? In this section, we seek to build an intuitive picture of the algorithms learned by two models on the discrimination task: CLIP ViT-B/16, and a randomly initialized ViT-B/16 (From Scratch).
[…]
Based on these attention patterns, we visualize how CLIP processes an image to solve the discrimination task in Figure 22. 1. Embedding: The model first tokenizes the image and embeds these tokens. Each object occupies four ViT-B/16 patches, so the objects are divided up into four tokens each. 2. Layer 1, Head 5: During the early perceptual stage, the local attention heads appear to aid in the formation of object representations by performing low-level comparisons within objects. For example, head 5 in layer 1 compares object tokens from left to right within each object. Other attention heads in this layer perform such comparisons in other directions, such as right to left or top to bottom. 3. Layer 5, Head 9: Towards the end of the perceptual stage, all object tokens within a single object attend to all other tokens within the same object. The four object tokens comprising each object have been pushed together in the latent space, and the model now “sees” a single object as a whole. 4. Layer 6, Head 11: The model switches from predominantly local to predominantly global attention in this layer, and within-pair (WP) attention peaks. The whole-object representations formed during the perceptual stage now attend to each other, indicating that the model is comparing them. 5. Layer 10, Head 6: The model appears to utilize object tokens (and background tokens) to store information, possibly the classification decision).
Figure 21: Example attention head patterns for models trained on the discrimination task. (a) CLIP ViT-B/16: On the left is an example input image, which is fed into the model. The heatmap is the same as Figure 2a—the $x$ and $y$-axes denote model layer and head index respectively, and the colors indicate the type of attention head as defined in Section 3 (local heads$=$blue, global heads$=$red). The attention patterns of four attention heads for this input image are displayed in black & white heatmaps below; white indicates higher attention values. The src axis indicates the source token, which is visually marked—recall that each object occupies four tokens each. The dest axis indicates the destination token. Individual objects attend to themselves during the perceptual stage (layers 0-5); objects begin to attend to the other object during the relational stage (layer 6 onwards). (b) From Scratch ViT-B/16: The analysis in (a) is repeated for a from scratch model trained on discrimination. The attention patterns are less interpretable throughout.
Figure 22: How CLIP ViT-B/16 processes an example from the discrimination task. Four attention heads are randomly selected from different stages in CLIP and analyzed on a single input image (see Figure 21). 1. Embedding: The model first tokenizes the input image. Each object occupies four ViT patches. 2. Layer 1, Head 5: During the perceptual stage, the model first performs low-level visual operations between tokens of individual objects. This particular attention head performs left-to-right attention within objects. 3. Layer 5, Head 9: Near the end of the perceptual stage, whole-object representations have been formed. 4. Layer 6, Head 11: During the relational stage, the whole-object representations are compared. 5. Layer 10, Head 6: Object and background tokens are used as registers to store information—presumably the classification.
Appendix G Distributed Alignment Search Technical Details
@@
Appendix H Perceptual Stage Analysis Controls
@@
Appendix I Perceptual Stage Analysis: Other Models
See Figures 25, 26, 27, and 28 for DINO, ImageNet, MAE, and From Scratch DAS results. We see that models broadly exhibit less disentanglement than CLIP and DINOv2.
@@ Appendix I Perceptual Stage Analysis: Other Models
Appendix J Novel Representations Analysis Technical Details
@@
Appendix K Relational Stage Analysis: Further Results
K.1 CLIP B/16 RMTS Novel Representation Analysis
See Figure 29 for novel representation analysis on CLIP B/16, finetuned for the relational match to sample task.
@@
K.2 Novel Representation Analysis: Other Models
See Figures 30, 31, 32, 33, 34, 35, 36 for DINO Discrimination/RMTS, ImageNet Discrimination/RMTS, MAE Discrimination/RMTS and From Scratch Discrimination Novel Representation Analysis results.
@@
Figure 34: Novel Representation Analysis for MAE ViT-B/16 (Disc.).
Figure 35: Novel Representation Analysis for MAE ViT-B/16 (RMTS).
@@
K.3 Abstract Representations of Same and Different
We run the linear probe and linear intervention analysis from Section 5 on DINO B/16, ImageNet B/16, and MAE B/16. We find that the intervention works much less well on these models than on DINOv2 B/16, CLIP B/16 or CLIP B/32. This indicates that these models are not using one abstract representation of same and one representation of different that is agnostic to perceptual qualities of the input image.
Additionally, we try to scale the directions that we are adding to the intermediate representations by 0.5 and 2.0 for DINO and ImageNet pretrained models, and find that neither of these versions of the intervention work much better well for either model. See Figure 37 and 38 for DINO and ImageNet results. See Figure 39 for MAE model results.
@@
Figure 39: Linear probe & intervention analysis for MAE ViT-B/16.
Appendix L Pipeline Loss Technical Details
To instill two-stage processing in ViTs trained from scratch on discrimination, we shape the model’s attention patterns in different ways at different layers. In particular, within-object attention is encouraged in layers 3, 4, and 5, while between-object attention is encouraged in layers 6 and 7 (roughly following CLIP’s two-stage processing; see Figure 2a). For models trained on RMTS, we additionally encourage between-pair attention in layers 8 and 9 (see Figure 2c). Finally, whenever we add the disentanglement loss, it is computed in layer 3 only.
To encourage a particular type of attention pattern in a given layer, we first compute the attention head scores (according to Section 3) for a randomly selected subset of either 4, 6, or 8 attention heads in that layer. [^9] This selection is kept constant throughout training; i.e., the attention heads that receive the pipeline loss signal are randomly chosen before training but do not change throughout training. These scores are then averaged across the layer. The average attention head score is subtracted from $1$, which is the maximum possible score for a given attention type (i.e. WO, WP, and BP following Figure 2). This difference averaged across model layers is the pipeline loss term. In the case of within-object attention, an average score of $1$ means that each attention head in the selected subset only attends between object tokens within the same object; no other tokens attend to each other. Thus, using the difference between $1$ and the current attention head scores as the loss signal encourages the attention heads to assign stronger attention between tokens within objects and weaker attention between all other tokens. The same holds for WP and BP attention. However, the particular forms of the attention patterns are not constrained; for example, in order to maximize the WO attention score in a given layer, models could learn to assign 100% of their attention between two object tokens only (instead of between all four tokens), or from a single object token to itself. This flexibility is inspired by the analysis in Figure 21, which finds that within-object attention patterns can take many different configurations that might serve different purposes in the formation of object representations. The same is true for WP and BP patterns.
Appendix M Auxiliary Loss Ablations
In this section, we present ablations of the different auxiliary loss functions presented in Section 7. Notably, the pipeline loss consists of two or three modular components, depending on the task. These components correspond to the processing stages that they attempt to induce — within-object processing (WO), within-pair processing (WP), and between-pair processing (BP). For discrimination, we find that either WO or WP losses confer a benefit, but that including both results in the best performance.
For RMTS, we find that including all loss functions once again confers the greatest performance benefit. Notably, we find that ablating either the disentanglement loss or the WP loss completely destroys RMTS performance, whereas ablating WO loss results in a fairly minor drop in performance.
WO Loss WP Loss Train Acc. Test Acc. Comp. Acc.
77.3% 76.5% 75.9%
91.6% 88.8% 84.5%
93.0% 91.1% 87.5%
95.6%  (+18.3) 93.9%  (+17.4) 92.3%  (+16.4)
Table 7: Performance of ViT-B/16 trained from scratch on the discrimination task with auxiliary losses.
Disent. Loss WO Loss WP Loss BP Loss Test Acc. Comp. Acc.
50.1% 50.0%
49.4% 50.7%
49.3% 51.3%
50.0% 51.3%
48.9% 50.0%
85.6% 68.7%
50.0% 50.1%
51.0% 51.2%
61.3% 50.8%
83.2% 68.8%
49.8% 50.9%
49.9% 50.5%
51% 50.8%
87.7% 76.5%
50.1% 50.1%
91.4% 77.4%
Table 8: Performance of ViTs trained from scratch on RMTS with auxiliary losses.
Appendix N Compute Resources